the sum of four consecutive integers is 238. find the integers

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

the sum of four consecutive integers is 238. find the integers
-------
1st: x-2
2nd: x-1
3rd: x
4th: x+1
-------------
Sum = 4x-2 = 238
4x = 240
x = 60
---
1st: x-2 = 58
2nd: x-1 = 59
3rd; x = 60
4th: x+1 = 61
==============
Cheers,
Stan H.

Answer by Alan3354(69443)

the sum of four consecutive integers is 238. find the integers
--------
238/4 = 59.5, the center
--> 58,59,60,61